Find all radian solutions to -2cos2(theta) -5sin(theta) +3=0 where 0<= theta <= 2pi
Answer by nshah11(47)

Let theta = t for simplicity.
-2cos(2t) - 5sin(t) + 3 = 0
-2(1 - 2sin^2(t)) - 5sin(t) + 3 = 0
4sin^2(t) - 5sin(t) + 1 = 0
4sin^2(t) - 4sin(t) - sin(t) + 1 = 0
(4sin(t) - 1)(sin(t) - 1) = 0
t = 0.253, 2.89, pi/2